A carrom board (4ft each side) has the queen at the centre .The queen,hit by the striker moves to the front
edge,rebounds and goes in the hole behind the striking line .find the magnitude of displacement of the queen (a)from the centre to the front edge ,(b)from the front edge to the hole,(c)from the centre to the hole?
1
Expert's answer
2012-06-14T08:07:00-0400
(a)from the centre to the front edge
d = l/2 = 2 ft (is equal to the half
length of the edge of carrom)
(b)from the front edge to the hole
d =
sqrt(4^2 + 2^2) = sqrt(20) = 4.47 ft
(c)from the centre to the hole?
d =
sqrt(2^2+2^2) = sqrt(8) = 2.82 ft